# GeoLRM: Geometry-Aware Large Reconstruction Model for High-Quality 3D Gaussian Generation

### v1

@@ Abstract
In this work, we introduce the Geometry-Aware Large Reconstruction Model (GeoLRM), an approach which can predict high-quality assets with 512k Gaussians and 21 input images in only 11 GB GPU memory. Previous works neglect the inherent sparsity of 3D structure and do not utilize explicit geometric relationships between 3D and 2D images. This limits these methods to a low-resolution representation and makes it difficult to scale up to the dense views for better quality. GeoLRM tackles these issues by incorporating a novel 3D-aware transformer structure that directly processes 3D points and uses deformable cross-attention mechanisms to effectively integrate image features into 3D representations. We implement this solution through a two-stage pipeline: initially, a lightweight proposal network generates a sparse set of 3D anchor points from the posed image inputs; subsequently, a specialized reconstruction transformer refines the geometry and retrieves textural details. Extensive experimental results demonstrate that GeoLRM significantly outperforms existing models, especially for dense view inputs. We also demonstrate the practical applicability of our model with 3D generation tasks, showcasing its versatility and potential for broader adoption in real-world applications.
@@ 1 Introduction
To address these challenges, we introduce the geometry-aware large reconstruction model (GeoLRM) for 3D Guassian generation. Our method centres on a 3D-aware reconstruction transformer that eschews conventional representations like triplanes or pixel-aligned Gaussians in favour of a direct interaction within the 3D space. However, directly generating 3D Gaussians in the whole 3D space requires huge memory cost. To this end, we first propose a specialized proposal network to predict an occupancy grid from input images. Only the occupied voxels will be further processed to generate 3D Gaussian features. The proposed transformer replaces the dense cross attention with deformable cross attention [85]. By projecting the input 3D tokens onto the corresponding image planes, these tokens only focus on the most relevant features, which greatly improves the effectiveness.
@@ 2.2 Large Reconstruction Model
Different with optimization-based 3D reconstruction methods, large reconstruction models [18, 22, 54, 73, 65, 81, 62, 64] are able to reconstruct 3D shapes in a feed-forward way. As the pioneer work of this area, the LRM [18] illustrates that the transformer backbone can effectively leverage the power of large-scale datasets and translate image tokens into implicit 3D triplanes under multi-view supervision. Beyond LRM, Instant3D  [22] improves reconstruction quality with spare-view inputs. It employs a two-stage paradigm, which first generates four views with the diffusion model and then regresses NeRF [33] from generated multi-view images. Instead of NeRF, InstantMesh [71] utilizes mesh representation to reconstruct 3D objects, which adopts a differentiable iso-surface extraction module. However, many of works [54, 81, 73, 70] choose 3D Gaussians [21] as the outputs. GRM [73] proposes a transformer network to translate pixels to the set of pixel-aligned 3D Gaussians while LGM [54] uses an asymmetric UNet to predict and fuse 3D Gaussians. Compared with these methods, our GeoLRM projects multi-view features to the 3D space with cross-view attention mechanisms, which explicitly explores geometric knowledge.
@@
Early methods [6, 7, 15, 35, 51, 72, 37] in 3D generation area utilize 3D GANs to generate 3D-aware contents. Despite that some methods  [32, 32, 84, 30, 10, 49, 79] replace 3D GANs with 3D diffusion models for high-quality generation, their generalization ability is bounded by the limited training data. Recently, proposed in DreamFusion [40], score distillation sampling (SDS) requires no 3D data and is able to leverage the great power of 2D text-to-image diffusion models [44, 43, 42]. Specifically, it optimizes a randomly-initialized 3D model and diffuses the render images with a pretrained diffusion model. As the follow-up works [63, 9, 26, 61, 55, 75, 27, 77, 25, 23, 41], many methods have been proposed to accelerate the optimization process or improve 3D generation quality. Different with SDS-based methods, Zero-1-to-3 [28] finetunes the 2D diffusion models on a large-scale synthetic dataset to change the camera viewpoint of a given image. Similar to Zero-1-to-3, many other works [47, 60, 48, 74, 29, 67, 31, 69] aim to synthesize multi-view consistent images. Our method can reconstruct 3D contents based on these synthesis multi-view images.
@@ 3 Methodology
Figure 2: Pipeline of the proposed GeoLRM, a geometry-powered method for efficient images to 3D reconstruction. The process begins with the transformation of dense tokens into an occupancy grid via a Proposal Transformer, which captures spatial occupancy from hierarchical image features extracted using a combination of a convolutional layer and DINOv2 [38]. Sparse tokens representing occupied voxels are further processed through a Reconstruction Transformer that employs self-attention and deformable cross-attention mechanisms to refine geometry and retrieve texture details with 3D to 2D projection. Finally, the refined 3D tokens are converted into 3D Gaussians for real-time rendering.
@@ 3.1 Overview
Figure 2 illustrates the pipeline of our proposed method. Our approach takes a set of images $\{{I}^{i}\}_{i=1}^{N}$ with their corresponding intrinsic $\{{K}^{i}\}_{i=1}^{N}$ and extrinsic $\{{T}^{i}\}_{i=1}^{N}$ as input. Initially, a proposal transformer predicts an occupancy grid. Each occupied voxel within this grid is considered a 3D anchor point. These 3D anchor points are then processed by a reconstruction transformer, refining their geometry and retrieving textural details. The proposal and reconstruction transformers share the same model architecture, which is further discussed in Section 3.2. The outputs of the reconstruction transformer are decoded into Gaussian features with a shallow MLP for rendering. Loss functions are described in Section 3.3.
@@ 3.2 Model Architecture
We present a geometry-aware transformer architecture featuring a hierarchical image encoder for extracting high and low-level image feature maps, and an anchor point decoder for transforming these features into 3D representations.
@@
Anchor Point Decoder The anchor point decoder aims to efficiently lift image features to 3D. Previous methods [18, 54, 73, 65, 81] uses tri-planes or pixel-aligned Gaussians to represent 3D contents. However, these data structures make it hard to utilize the projection relationships, causing dense computations. Instead, we use 3D anchor points, which serve as proxies for their surrounding points, significantly reducing the number of points we need to process. As detailed in Figure 2, each decoder block contains a self-attention layer, a deformable cross-attention layer and a feed-forward network (FFN). The model takes $N$ anchor point features $\mathcal{F}_{A}=\{\boldsymbol{f}_{i}\}_{i=1}^{N}$ as input tokens. Each token $\boldsymbol{f}_{i}$ comprises the coordinate of the corresponding point and a shared learnable feature.
@@
Given input tokens $\mathcal{F}_{A}^{in}$, the decoder block enhances these tokens through a series of sophisticated transformations:
@@
Post Processing The proposal network takes a low-resolution dense grid ($16^{3}$) as anchor points. The output is upsampled to a high-resolution grid ($128^{3}$) with a linear layer. This grid is formulated to represent the occupancy probability of the corresponding area ($[-0.5,0.5]^{3}$). The reconstruction transformer takes occupied voxels as anchor points. Each output token $\boldsymbol{f}_{i}$ is decoded into multiple 3D Gaussians $\{\boldsymbol{G}_{ij}\}_{j=1}^{M}$ with a linear layer. The 3D Gaussian $\boldsymbol{G}_{ij}$ is parameterized by the offset $\boldsymbol{o}_{ij}$ regarding the anchor points, 3-channel RGB $\boldsymbol{c}_{ij}$, 3-channel scale $\boldsymbol{s}_{ij}$, 4-channel rotation quaternion $\boldsymbol{\sigma}_{ij}$, and 1-channel opacity $\alpha_{ij}$. We employ activation functions to limit the range of the offset, scale and opacity for better training stability similar to [54]:
@@ 3.3 Training Objectives
We employ a two-stage training mechanism for our model. In the first stage, we train the proposal transformer using 3D occupancy ground truth. This stage presents a challenge as it involves a highly unbalanced binary classification task; only about 5% of the voxels are occupied. To address this imbalance, we employ a combination of binary cross-entropy loss and the scene-class affinity loss, as proposed in [5], to supervise the training process. For the generation of ground truth data, see A.2.
@@ 4.1 Datasets
Google Scanned Objects (GSO) [13]: Used for evaluation, this dataset is rendered similarly to GObjaverse to maintain consistency. We randomly select a subset of 100 objects to streamline the evaluation process.
4.2 Implimentation details
Our model features 330 million parameters distributed across two distinct image encoders and two transformers. The first encoder processes geometry with the 6-layer proposal transformer, while the second focuses more on textures crucial with the 16-layer reconstruction transformer. During training, we maintain a maximum number of transformer input tokens of 4k and randomly select 8 views from a possible 38 for supervision. From these 8 views, we randomly select 1 to 7 views as inputs to predict the remaining views. This flexibility in view selection not only tests the robustness of our method but also mimics real-world scenarios where complete data may not always be available. Both input and rendering resolutions are maintained at 448x448 pixels. At the testing and inference stages, the model processes up to 16k input tokens, showcasing its scalability without the need for fine-tuning. Detailed information on our model’s architecture and training procedures can be found in Section A.3.
@@ 4.3 Quantitative Results
Table 1: Quantitative results on Google Scanned Objects (GSO) [13], where we used six views for inputs and four for evaluation. Bold and underline denote the highest and second-highest scores, respectively.
@@
LGM [54] 20.76 0.832 0.227 0.295 0.703
CRM [64] 22.78 0.843 0.190 0.213 0.831
InstantMesh [71] 23.19 0.856 0.166 0.186 0.854
Ours 23.42 0.865 0.174 0.165 0.890
Table 2: Quantitative results on Google Scanned Objects (GSO) [13] with different numbers of input views. We keep the same four views for testing while changing the number of input views. Bold denotes the highest score.
4 Inputs 8 Inputs 12 Inputs
Mothod PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
InstantMesh [71] 22.87 0.832 23.22 0.861 23.05 0.843
Ours 22.64 0.840 23.76 0.870 24.39 0.887
We evaluated the quality of reconstructed assets from sparse view inputs by analyzing both 2D visual and 3D geometric aspects on the GSO dataset [13]. Visual quality was assessed by comparing rendered views to ground truth images using metrics such as PSNR, SSIM, and LPIPS. Geometric accuracy was evaluated by aligning our models to the ground truth coordinate systems and measuring discrepancies using Chamfer Distance and F-Score at a threshold of 0.2, with point samples totalling 16,000 from the ground truth surfaces. Our method was quantitatively compared against established baselines, including LGM [54], CRM [64], and InstantMesh [71]. We avoided comparisons with proprietary methods due to the unavailability of their test splits. Similarly, we excluded comparisons with OpenLRM [17] and TripoSR [56] as these methods are tailored for single image inputs, which would be unfair to compare with.
Our approach achieved state-of-the-art performance in four out of the five metrics studied. Although InstantMesh showed slightly higher LPIPS, attributed to its mesh-based smoothing capabilities, our method demonstrated superior geometric accuracy, benefiting from explicit modelling of the 3D-to-2D relationship.
In another experiment, outlined in Table 2, we observed a notable trend: the performance of our model improves consistently as the number of input views increases. This indicates robust scalability, a critical feature for practical applications. In contrast, the performance of InstantMesh [71], does not follow this pattern. Specifically, InstantMesh shows a decline in performance when the input views increase to 12. This degradation could be due to two primary factors. First, the low-resolution tri-planes may reach their maximum capacity to represent details. Second, the model tends to oversmooth details when handling a large volume of image tokens. Our approach strategically addresses these issues. We employ an extendable sequence of 3D tokens that can be dynamically adjusted to fit the resolution requirements. Additionally, our model features deformable attention mechanisms that intelligently focus on the most pertinent information, preventing the loss of critical details.
@@ 4.4 Qualitative Results
We conducted a qualitative analysis comparing our method with several LRM-based baselines, including TripoSR [17], LGM [54], CRM [64], and InstantMesh [71], maintaining their original settings to ensure optimal performance. In our approach, we utilized the SV3D [60] technology to generate 21 multi-view images, significantly enhancing the resolution and textural details of the 3D Gaussians produced, as illustrated in Figure 3. Furthermore, as shown in Figure 4, employing InstantMesh to reconstruct these images did not yield satisfactory outcomes, corroborating our quantitative findings. This demonstrates the superior capability of our method in handling complex 3D reconstructions.

### v2

@@ Abstract
In this work, we introduce the Geometry-Aware Large Reconstruction Model (GeoLRM), an approach which can predict high-quality assets with 512k Gaussians and 21 input images in only 11 GB GPU memory. Previous works neglect the inherent sparsity of 3D structure and do not utilize explicit geometric relationships between 3D and 2D images. This limits these methods to a low-resolution representation and makes it difficult to scale up to the dense views for better quality. GeoLRM tackles these issues by incorporating a novel 3D-aware transformer structure that directly processes 3D points and uses deformable cross-attention mechanisms to effectively integrate image features into 3D representations. We implement this solution through a two-stage pipeline: initially, a lightweight proposal network generates a sparse set of 3D anchor points from the posed image inputs; subsequently, a specialized reconstruction transformer refines the geometry and retrieves textural details. Extensive experimental results demonstrate that GeoLRM significantly outperforms existing models, especially for dense view inputs. We also demonstrate the practical applicability of our model with 3D generation tasks, showcasing its versatility and potential for broader adoption in real-world applications. The project page: https://linshan-bin.github.io/GeoLRM/.
@@ 1 Introduction
To address these challenges, we introduce the geometry-aware large reconstruction model (GeoLRM) for 3D Gaussian generation. Our method centres on a 3D-aware reconstruction transformer that eschews conventional representations like triplanes or pixel-aligned Gaussians in favour of a direct interaction within the 3D space. However, directly generating 3D Gaussians in the whole 3D space requires huge memory costs. To this end, we first propose a specialized proposal network to predict an occupancy grid from input images. Only the occupied voxels will be further processed to generate 3D Gaussian features. The proposed transformer replaces the dense cross attention with deformable cross attention [86]. By projecting the input 3D tokens onto the corresponding image planes, these tokens only focus on the most relevant features, which greatly improves the effectiveness.
@@
In summary, our contributions are as follows:
  • We introduce a two-stage pipeline that leverages the sparse nature of 3D data, resulting in a sparse 3DGS token representation suitable for extension to high resolution.
  • We fully exploit the projection relationship between 3D points and 2D images, significantly reducing the space complexity of attention mechanisms in LRMs, thus enabling denser image input configurations.
  • To the best of our knowledge, GeoLRM is the first to process dense inputs using LRM, potentially paving the way for integrating video generation models into 3D AIGC applications.
@@ 2.2 Large Reconstruction Model
Different from optimization-based 3D reconstruction methods, large reconstruction models [18, 22, 54, 74, 65, 82, 62, 64] are able to reconstruct 3D shapes in a feed-forward way. As the pioneer work of this area, the LRM [18] illustrates that the transformer backbone can effectively leverage the power of large-scale datasets and translate image tokens into implicit 3D triplanes under multi-view supervision. Beyond LRM, Instant3D  [22] improves reconstruction quality with sparse-view inputs. It employs a two-stage paradigm, which first generates four views with the diffusion model and then regresses NeRF [33] from generated multi-view images. Instead of NeRF, InstantMesh [72] utilizes mesh representation to reconstruct 3D objects, which adopts a differentiable iso-surface extraction module. However, many works [54, 82, 74, 71] choose 3D Gaussians [21] as the outputs. GRM [74] proposes a transformer network to translate pixels to the set of pixel-aligned 3D Gaussians while LGM [54] uses an asymmetric UNet to predict and fuse 3D Gaussians. Compared with these methods, our GeoLRM projects multi-view features to the 3D space with cross-view attention mechanisms, which explicitly explores geometric knowledge.
@@
Early methods [6, 7, 15, 35, 51, 73, 37] in 3D generation area utilize 3D GANs to generate 3D-aware contents. Although some methods  [32, 32, 85, 30, 10, 49, 80] replace 3D GANs with 3D diffusion models for high-quality generation, their generalization ability is bounded by the limited training data. Recently, proposed in DreamFusion [40], score distillation sampling (SDS) requires no 3D data and is able to leverage the great power of 2D text-to-image diffusion models [44, 43, 42]. Specifically, it optimizes a randomly-initialized 3D model and diffuses the render images with a pretrained diffusion model. As the follow-up works [63, 9, 26, 61, 55, 76, 27, 78, 25, 23, 41], many methods have been proposed to accelerate the optimization process or improve 3D generation quality. Different with SDS-based methods, Zero-1-to-3 [28] fine-tunes the 2D diffusion models on a large-scale synthetic dataset to change the camera viewpoint of a given image. Similar to Zero-1-to-3, many other works [47, 60, 48, 75, 29, 67, 31, 69] aim to synthesize multi-view consistent images. Our method can reconstruct 3D contents based on these synthesis multi-view images.
@@ 3 Methodology
Figure 2: Pipeline of the proposed GeoLRM, a geometry-powered method for efficient image to 3D reconstruction. The process begins with the transformation of dense tokens into an occupancy grid via a Proposal Transformer, which captures spatial occupancy from hierarchical image features extracted using a combination of a convolutional layer and DINOv2 [38]. Sparse tokens representing occupied voxels are further processed through a Reconstruction Transformer that employs self-attention and deformable cross-attention mechanisms to refine geometry and retrieve texture details with 3D to 2D projection. Finally, the refined 3D tokens are converted into 3D Gaussians for real-time rendering.
@@ 3.1 Overview
Figure 2 illustrates the pipeline of our proposed method. Our approach takes a set of images $\{{I}^{i}\}_{i=1}^{N}$ with their corresponding intrinsic $\{{K}^{i}\}_{i=1}^{N}$ and extrinsic $\{{T}^{i}\}_{i=1}^{N}$ as input. Initially, we encode input images into hierarchical image features and predict an occupancy grid with a proposal transformer. Each occupied voxel within this grid is considered a 3D anchor point. These 3D anchor points are then processed by a reconstruction transformer, refining their geometry and retrieving textural details. The proposal and reconstruction transformers share the same model architecture, which is further discussed in Section 3.2. The outputs of the reconstruction transformer are decoded into Gaussian features with a shallow MLP for rendering. Loss functions are described in Section 3.3.
@@ 3.2 Model Architecture
Our model architecture features a hierarchical image encoder for extracting high and low-level image feature maps along with a geometry-aware transformer for lifting 2D features into 3D representations.
@@
Geometry-aware Transformer The geometry-aware transformer aims to efficiently lift image features to 3D. The proposal transformer and reconstruction transformer are both instances of this architecture. Previous methods [18, 54, 74, 65, 82] use tri-planes or pixel-aligned Gaussians to represent 3D contents. However, these data structures make it hard to utilize the projection relationships, causing dense computations. Instead, we use 3D anchor points, which serve as proxies for their surrounding points, significantly reducing the number of points we need to process. As detailed in Figure 2, each transformer block contains a self-attention layer, a deformable cross-attention layer and a feed-forward network (FFN). The model takes $N$ anchor point features $\mathcal{F}_{A}=\{\boldsymbol{f}_{i}\}_{i=1}^{N}$ as input tokens. Each token $\boldsymbol{f}_{i}$ comprises the coordinate of the corresponding point and a shared learnable feature.
@@
Given input tokens $\mathcal{F}_{A}^{in}$, the transformer block enhances these tokens through a series of sophisticated transformations described as follows:
@@
Post-processing The proposal network takes a low-resolution dense grid ($16^{3}$) as anchor points. The output is upsampled to a high-resolution grid ($128^{3}$) with a linear layer. This grid is formulated to represent the occupancy probability of the corresponding area ($[-0.5,0.5]^{3}$). The reconstruction transformer takes occupied voxels as anchor points. Each output token $\boldsymbol{f}_{i}$ is decoded into multiple 3D Gaussians $\{\boldsymbol{G}_{ij}\}_{j=1}^{M}$ with a multilayer perceptron. The 3D Gaussian $\boldsymbol{G}_{ij}$ is parameterized by the offset $\boldsymbol{o}_{ij}$ regarding the anchor points, 3-channel RGB $\boldsymbol{c}_{ij}$, 3-channel scale $\boldsymbol{s}_{ij}$, 4-channel rotation quaternion $\boldsymbol{\sigma}_{ij}$, and 1-channel opacity $\alpha_{ij}$. We employ activation functions to limit the range of the offset, scale and opacity for better training stability similar to [54]:
@@ 3.3 Training Objectives
We employ a two-stage training mechanism for our model. In the first stage, we train the proposal transformer using 3D occupancy ground truth. This stage presents a challenge as it involves a highly unbalanced binary classification task; only about 5% of the voxels are occupied. To address this imbalance, we employ a combination of binary cross-entropy loss and the scene-class affinity loss, as proposed in [5], to supervise the training process. For the generation of ground truth data, see A.1.
@@ 4.1 Datasets
Google Scanned Objects (GSO) [13]: Used for evaluation, this dataset is rendered similarly to GObjaverse. A random subset of 100 objects is selected to streamline the evaluation process.
OmniObject3D [70]: Also used for evaluation, this dataset is consistently rendered like GObjaverse. A random subset of 100 objects is chosen for efficient evaluation.
4.2 Implementation details
Our model features 330 million parameters distributed across two distinct image encoders and two transformers. The first encoder processes geometry with the 6-layer proposal transformer, while the second focuses more on textures crucial with the 16-layer reconstruction transformer. During training, we maintain a maximum number of transformer input tokens of 4k and randomly select 8 views from a possible 38 for supervision. From these 8 views, we randomly select 1 to 7 views as inputs to predict the remaining views. This flexibility in view selection not only tests the robustness of our method but also mimics real-world scenarios where complete data may not always be available. Both input and rendering resolutions are maintained at 448x448 pixels. At the testing and inference stages, we use a resolution of 512x512 to align with existing methods. Besides, the number of input tokens is extended to 16k during testing, showcasing its scalability without the need for fine-tuning. Detailed information on our model’s architecture and training procedures can be found in Section A.3.
@@ 4.3 Quantitative Results
Table 1: Quantitative results on Google Scanned Objects (GSO) [13], where we used six views for inputs and four for evaluation. Inference time and memory usage account only for the reconstruction process. Bold and underline denote the highest and second-highest scores, respectively.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ CD $\downarrow$ FS $\uparrow$ Inf. Time (s) Memory (GB)
LGM 20.76 0.832 0.227 0.295 0.703 0.07 7.23
CRM 22.78 0.843 0.190 0.213 0.831 0.30 5.93
InstantMesh 23.19 0.856 0.166 0.186 0.854 0.78 23.12
Ours 23.57 0.872 0.167 0.167 0.892 0.67 4.92
Table 2: Quantitative results on OmniObject3D [70]. Bold and underline denote the highest and second-highest scores, respectively.
@@
LGM 21.94 0.824 0.203 0.256 0.787
CRM 23.12 0.855 0.175 0.204 0.810
InstantMesh 23.86 0.860 0.139 0.178 0.834
Ours 24.74 0.883 0.134 0.156 0.863
Table 3: Quantitative results on Google Scanned Objects (GSO) with different numbers of input views. We keep the same four views for testing while changing the number of input views. Bold denotes the highest score.
Num Input PSNR SSIM Inf. Time (s) Memory (GB) InstantMesh Ours InstantMesh Ours InstantMesh Ours InstantMesh Ours 4 22.87 22.84 0.832 0.851 0.68 0.51 22.09 4.30 8 23.22 23.82 0.861 0.883 0.87 0.84 24.35 5.50 12 23.05 24.43 0.843 0.892 1.07 1.16 24.62 6.96 16 23.15 24.79 0.861 0.903 1.30 1.51 26.69 8.23 20 23.25 25.13 0.895 0.905 1.62 1.84 28.73 9.43
We evaluated the quality of reconstructed assets from sparse view inputs by analyzing both 2D visual and 3D geometric aspects on the GSO and OmniObject3D dataset [13]. Visual quality was assessed by comparing rendered views to ground truth images using metrics such as PSNR, SSIM, and LPIPS. Geometric accuracy was evaluated by aligning our models to the ground truth coordinate systems and measuring discrepancies using Chamfer Distance and F-Score at a threshold of 0.2, with point samples totalling 16,000 from the ground truth surfaces. Our method was quantitatively compared against established baselines, including LGM [54], CRM [64], and InstantMesh [72]. We avoided comparisons with proprietary methods due to the unavailability of their test splits. Similarly, we excluded comparisons with OpenLRM [17] and TripoSR [56] as these methods are tailored for single image inputs, which would be unfair to compare with.
Our approach achieved state-of-the-art performance in four out of the five metrics studied. Although InstantMesh showed slightly higher LPIPS on the GSO dataset, attributed to its mesh-based smoothing capabilities, our method demonstrated superior geometric accuracy, benefiting from explicit modelling of the 3D-to-2D relationship.
In another experiment, outlined in Table 3, we observed a notable trend: our model’s performance consistently improves with more input views while maintaining low computational costs. This indicates robust scalability, a critical feature for practical applications. In contrast, the performance of InstantMesh [72], does not follow this pattern. Specifically, InstantMesh shows a decline in performance when the input views increase to 12. This degradation could be due to two primary factors. First, the low-resolution tri-planes may reach their maximum capacity to represent details. Second, the model tends to oversmooth details when handling a large volume of image tokens. Our approach strategically addresses these issues. We employ an extendable sequence of 3D tokens that can be dynamically adjusted to fit the resolution requirements. Additionally, our model features deformable attention mechanisms that intelligently focus on the most pertinent information, preventing the loss of critical details.
@@ 4.4 Qualitative Results
We conducted a qualitative analysis comparing our method with several LRM-based baselines, including TripoSR [17], LGM [54], CRM [64], and InstantMesh [72], maintaining their original settings to ensure optimal performance. In our approach, we utilized the SV3D [60] technology to generate 21 multi-view images, significantly enhancing the resolution and textural details of the 3D Gaussians produced, as illustrated in Figure 3. Furthermore, as shown in Figure 4, employing InstantMesh to reconstruct these images did not yield satisfactory outcomes, corroborating our quantitative findings. This demonstrates the superior capability of our method in handling more complex 3D reconstructions.